Suppose that the risk-free rate is Rf = 2.53% and the risk-premium is E(Rm) − Rf = 7.10%. According to Gordon’s Growth Model, if a company has a current dividend of D0 = $22.33 per share, a constant growth rate of g = 5.24%, and β = 1.21, what is its stock price?
Answer #1
As per CAPM Model,
Er = Rf + (Rm-Rf)*B
Where,
Er= Required return
Rf= Risk free return
ERm= Market return
B= Beta
Er = 2.53 + (7.10*1.21)
Er = 2.53 + 8.591
Er = 11.121%

Therefore, required return = 11.121%

Value of the company now using Dividend Growth model
In dividend growth model,
P = D0(1+g)/Er-g
Where,
P= price of the share
D0= Current dividend
g= growth rate
Er= required return
Current value of the company,
P = $22.33(1.0524)/(0.11121-0.0524)
P = $23.5 / 0.5881
P = $399.59

Therefore, required stock price = $399.59